Annette drive to Shawnee in 4 hours and drove back in three hours. What were her speeds of her speed coming back was 11 miles per hour greater that her speed going?
Answer by macston(5194)

.
D=distance one way; R=Annette's speed going to Shawnee
R+11mph=speed returning
.
4(R)=3(R+11mph)
4R=3R+33mph
R=33mph
ANSWER 1: Her speed going was 33 mph.
.
Return speed=R+11mph=33mph+11mph=44mph
ANSWER 2: Her return speed was 44 mph.
.
CHECK:
D=R(4hrs)
D=33mph(4hrs)=132 miles
.
D=R+11(3hrs)
132mi=(33mph+11mph)(3hrs)
132mi=(44mph)(3hrs)
132mi=132mi